Show your calculation and explain how the base 60 number system is used to subtract 48 minutes from 2 hour 13 minutes.

My calculation (3)

Explanation of my steps (3)

 

 


1
Expert's answer
2022-05-23T16:48:38-0400
"\begin{matrix}
 & 2h &13min \\
 - \\
 & & 48min
\end{matrix}"

"1h=60min"



"\begin{matrix}
 & (2-1)h &(13+60)min \\
 - \\
 & & 48min
\end{matrix}"



"\begin{matrix}
 & 1h &73min \\
 - \\
 & & 48min \\
\hdashline \\
 & 1h &25min \\
\end{matrix}"